# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v6

@@ Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. This paper first presents a data structure, GLOVER (Graph Level Order), which is a partitioning of the nodes into a total order of neighborhoods. This data structure allows us to introduce a total order on oriented graphs and construct a decreasing sequence of subsets. This decreasing sequence of subsets is what ensures that no counterexample can exist. Further, this algorithmic approach lets us know where the node is. Later findings extend the results to multiple nodes. This is true for all oriented graphs. These findings also have practical applications in network optimization and algorithm design.
@@ 1 Introduction
Seymour’s Second Neighborhood Conjecture (SSNC) was originally proposed by Paul Seymour in 1990. It asks the question of the existence of a node in an oriented graph whose out-degree at least doubles in the graph’s square. The conjecture was first published by Nathaniel Dean and Brenda J. Latka in 1995 [4], where they posed a similar conjecture specific to tournaments. This version of the conjecture was first proven by Fischer [8], with a second proof provided by Havet and Thomasse [10].
@@
(Seymour’s Second Neighborhood Conjecture). For every oriented graph $G=(V,E)$, there exists a vertex $v\in V$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}$ represent the first and second neighborhoods of $v$ respectively.
This work generalizes an approach initially developed for the minimum degree 7 case (outlined in Glover [9]). Here, we hope to tackle the entire class of oriented graphs based on the structure laid out in that paper.
@@
The Seymour Second Neighborhood Challenge is a difficult graph theory problem. The problem itself asks us to focus on the question of oriented graphs and second neighbors. There have been several papers that have looked into special cases. Kaneko and Locke [11] showed that the conjecture was true for graphs with a minimum degree of at most six. Brantner et al. [1] showed that, if the conjecture is false, infinitely many counterexamples must exist. Additionally, Chen et al. [3] provided a lower bound for how much a node’s out-degree increases in the graph’s square. They concluded that every oriented graph has a vertex whose second neighborhood is at least $\lambda$ times larger than the first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1. In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Bolter et al. [2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the Seymour Second Neighborhood Conjecture.
The topic of $m$-free graphs is of particular interest to this paper because Daamouch’s [7] work is very similar to what we see here. In his paper, he introduces induced subgraphs and speaks of their roles in degree properties on vertices, which will also be a key thing in this paper.
The Seymour Second Neighborhood Conjecture is not strictly a graph theory problem. Although it has historically been studied within the field of mathematics, it has practical applications across several domains, including social media, social networking, epidemiology, disease control, and many other fields. In the critical context of disease control, the nodes in this conjecture can represent individuals, and the orientation of the arcs symbolizes one person infecting another. Here, the application of SSNC becomes relevant by asking: Is there a person whose influence in their second radius of interaction (i.e., people they indirectly connect with) exceeds the number of people in their first radius (i.e., direct contacts)?
In this paper, we will prove the Seymour Second Neighborhood Conjecture is true. We will do this by first assuming that it is false and attempting to construct a counterexample. We will explore the conditions under which no node’s out-degree doubles will double in the square of the oriented graph. This approach will focus initially on minimum out-degree nodes. We will analyze how their out-degree could fail to double. We will then extend our algorithm to other nodes based on their distance from the minimum out-degree node.
The paper is organized as follows: Section 2 begins with some common terminology used in graph theory. Section 3 introduces our approach to the problem. In section 4 we look at some small examples of this style of thinking. Section 5 introduces the terminology and identifies the conditions under which their out-degree might fail to double. Section 6 examines the back arcs and their place in oriented graphs. Section 7 looks at the conditions that would cause an interior degree to double and its importance to the Seymour Second Neighborhood Conjecture. Section 8 further looks at containers under the context of density constraints. In Section 9, we consider transitive triangles and Seymour diamonds and show that they are not a hindrance to our algorithm. Section 10 presents this process, the full Decreasing Sequence Property algorithm, as well as the theorem stating its solution to the Conjecture.
@@ 2 Graph Theory Terminology
Now that we have two different conjectures, we can investigate how they relate to one another. Conjecture 1.1 relates the notion of a node’s degree doubling to its neighborhoods, while conjecture 2.1 relates this same concept to the square graph. The defining characteristics of these conjectures are the relationship between a node’s out-degree at least doubling, and its second neighborhood being greater than or equal to the first. By the definition of an arc in the square graph, we can see that an arc is defined in this square graph only for paths of length two in the original graph. This is why these two conjectures are seen as identical.
We will focus our attention on Conjecture 1.1 for the reason that the conjecture 2.1 requires two different graphs - the original graph and the square graph. While the construction of the square graph is possible, we have found that it is not as helpful to our research at this time. Instead, focusing on the neighborhood analysis brought many insights and characteristics that helped to develop our algorithm.
Next, we turn our attention to the term degree-doubling nodes. By this term, we aim to get to the heart of conjecture 1.1. That conjecture is looking for a node whose degree at least doubles, and we are doing so by looking at the first and second neighborhoods of nodes and finding one whose second neighborhood is at least as large as the first, as the conjecture asks for. A similar definition, a Seymour vertex, fulfills conjecture 2.1 - it’s out-degree in the squared graph is at least double its original out-degree. We could equivalently search for a Seymour vertex or a degree-doubling node.
@@ 3 Approach and Motivation
From this foundation, we can now partition the graph into $d_{v_{0}}+1$ subgraphs corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$. This partitioning enables us to leverage the divide-and-conquer paradigm. No nodes are being removed to form these partitions. This is a conceptual idea. Instead, think of it as putting a box around the partition. Now the nodes inside the box can work with the other nodes inside the box and solve their problem independent of the other partitions. These partitions are still connected to one another, but thinking of them in this way allows us to do more.
This represents a stark contrast to most previous approaches to the Seymour Second Neighborhood Conjecture. Many of these previous approaches were conceptual or theoretical. They would say only that a node whose degree doubled existed, but the notion of a path to that node was not part of the paper. Secondly, most of these papers focused on local search techniques that started at or near the desired degree-doubling node. That presents a problem because oriented graphs can be varying in nature. This is the first exhaustive search technique on the SSNC problem.
When speaking of an exhaustive search algorithm, complexity becomes a factor and questions of how the algorithms will handle large graphs compared to smaller graphs. This should not be a problem as we will see in the following sections; the slowest part of this algorithm is building the data structure. That will require partitioning the nodes and edges of the graph, in a worst case complexity of $O(V+E)$.
By partitioning the graph, even larger graphs, often with higher minimum degree nodes and thus more partitions, are able to be broken into sub-problems and solved independent of one another. These subproblems are each their own instances of the Seymour Second Neighborhood Conjecture, just a localized version. What the partition is doing is breaking a node’s individual degree into two parts-the degree within the container and the degree to the next container. The degree within the container is the subproblem that will be solved separately, and the degree to the next container is how they connect back to the global problem.
We still need to show that this is a formidable foundation. We have laid the groundwork for our algorithm to tackle the oriented graphs. With a minimum degree node, we can partition this graph into independent subproblems. This is now ready for an exhaustive search algorithm.
@@ 4 Initial Lemmas
These first lemmas will seek to place a lower bound on the out-degree of a minimum out-degree node in a counterexample. For this section, unless otherwise noted, assume that $v_{0}$ is a node with minimum out-degree in our oriented graph $G$. We will also introduce some standard notation from graph theory. The work in this section agrees with what was already done by [11] and is not meant to serve as original content. It does, however, provide insight into our reasoning.
@@
Minimum Out-Degree $<$ 3 Suppose a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double in $G^{2}$.
@@
Case 1: Assume that $d^{+}(v_{0})=0$. Then $v_{0}$ has no outgoing edges in $G$. In $G^{2}$, it will still have out-degree 0. Hence its out-degree trivially satisfies the condition of at least doubling.
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. We know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. This node $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$. This means that $(v_{1},v_{2})\in G$. This implies that in $G^{2}$, both $(v_{0},v_{1})$ and $(v_{0},v_{2})$ exist, giving $v_{0}$ an out-degree of at least 2, thus doubling its out-degree.
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. We know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. Consider the induced subgraph $G[v_{1},v_{2}]$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. The node $v_{0}$ cannot be one of these nodes. Hence, in $G^{2}$, $v_{0}$ will connect to $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, giving it an out-degree of at least 4. Thus, it’s out-degree will double.
Therefore, in all cases, $v_{0}$’s out-degree at least doubles in $G^{2}$.
Lemma 4.1 Minimum Out-Degree $<$ 3 begins to tell us about the structure of oriented graphs. We are learning that, in order to construct our counterexample, we need our minimum degree to be at least 3. This will also place that same minimum value restriction on every other node in the graph. In each of the cases of the lemma, the minimum degree node has few very out-neighbors, which couldn’t stop the increase that was coming from the influx of second neighbors.
@@
Minimum Out-Degree 3 with Neighbor Out-Degree 0 Suppose a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3. Suppose also that at least one of $v_{0}$’s neighbors has out-degree 0; then the out-degree of $v_{0}$ will at least double in $G^{2}$.
@@
Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, and $v_{3}\in V(G)$. Consider the induced subgraph $G[v_{1},v_{2},v_{3}]$. Now, assume that $v_{1}$ has an out-degree of 0 within $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to 3 nodes outside of $\{v_{1},v_{2},v_{3}\}$.
Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$. As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in $G$.
@@
Figure 3: This is a representation of the node $v_{0}$ in $G$, $v_{0}$ being a minimum degree node, connecting to three nodes, $v_{1}$, $v_{2}$, and $v_{3}$. Each of these nodes ($v_{1}$, $v_{2}$, and $v_{3}$) connects to the nodes $v_{4}$ and $v_{5}$. The neighbor of $v_{4}$ is $v_{5}$, and the neighbors of $v_{5}$ are $v_{6}$, $v_{7}$, and $v_{8}$. While $v_{0}$ has degree 3, the neighbors of $v_{0}$ all have degree 1 in the subgraph $G[N^{+}(v_{0})]$, and they each have degree 2 outside $G[N^{+}(v_{0})]$. Thus, in $G^{2}$, the degree of $v_{0}$ will double.
Minimum Out-Degree 3 with Neighbors 1
@@
Assume that in an oriented graph, the minimal out-degree node $v_{0}$ has an out-degree of 3. Additionally, suppose that in the induced subgraph of its neighbors, every neighbor has an out-degree of 1. The out-degree of $v_{0}$ or one of its neighbors will then at least double in $G^{2}.$.
@@
Each node $v_{1},v_{2},v_{3}$ must connect to two nodes in $G-G[v_{1},v_{2},v_{3}].$ If the set of nodes to which $v_{1},v_{2},v_{3}$ connect in $G-G[v_{1},v_{2},v_{3}]$ contains more than 2 distinct nodes, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume they all connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
This reduces the situation to Case 3 in Lemma 4.1 Minimum Out-Degree $<$ 3, where $v_{1}$ acts similarly to how $v_{0}$ did in that case. There is at most one edge between $v_{4}$ and $v_{5}$. One of these nodes, say $v_{4}$, must connect to three additional nodes, $v_{6},v_{7}$, and $v_{8}$. As a result, in $G^{2}$, we will have the edges $(v_{1},v_{6})$, $(v_{1},v_{7})$, and $(v_{1},v_{8})$, meaning that the out-degree of $v_{1}$ will at least double in $G^{2}$. ∎
Lemma 4 (Minimum Out-Degree 3 with Neighbors 1) provides an exhaustive search strategy, which is a question we shall revisit throughout this paper. It must prevent a node’s out-degree from doubling in order for that to function. In this instance, we observe that the degree of the starting node in the example $v_{0}$ is not doubled. We are compelled to look for the next node instead. In this case, the degree of the neighbor of $v_{0}$ $v_{1}$ doubles.
We did not require the ”right” starting node; therefore, this is a crucial distinction. Not because of its possible proximity to the double degree node, but rather because of its capacity to divide the nodes in the graph, we selected a minimal degree node.
We shall use double arrows to show that a node or collection of nodes connects to every node in the next subgraph for graphs similar to those in Lemma 4.2 throughout the rest of this work. Because all of the nodes in a shared induced subgraph relate to the same set of nodes, this is possible.

### v7

@@ Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. In this paper, we first present a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers. This data structure establishes a well-ordering on oriented graphs and allows for the construction of a decreasing sequence of subsets of nodes. This sequence proves the non-existence of counterexamples to the SSNC and precisely identifies the required node. Further, our approach finds the occurrence of dense regular graphs inside containers. This finding extends the SSNC to the discovery of multiple nodes satisfying the degree doubling property. Beyond theoretical implications, the algorithm and data structure have practical applications in network optimization and algorithm design.
@@ 1 Introduction
The Seymour Second Neighborhood Conjecture (SSNC), proposed by Paul Seymour in 1990, posits that in every oriented graph, there exists a node whose out-degree in the graph’s square is at least twice its out-degree in the original graph. This conjecture was first formally published by Dean and Latka in 1995 [4], where they proposed a related version specific to tournaments. The tournament case was proved by Fisher [8], with an alternative proof later provided by Havet and Thomassé [10].
@@
(Seymour’s Second Neighborhood Conjecture). For every oriented graph $G=(V,E)$, there exists a vertex $v\in V$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}(v)$ denote the first and second neighborhoods of $v$, respectively.
@@
Despite decades of research, the SSNC remains one of graph theory’s challenging open problems. Special cases of the conjecture have been explored extensively. Kaneko and Locke [11] showed that the SSNC holds for graphs with a minimum degree of at most six. Brantner et al. [1] demonstrated that if the conjecture were false, infinitely many counterexamples would exist. Chen et al. [3] established a lower bound, proving that every oriented graph contains a vertex whose second neighborhood is at least $lambda$ times larger than its first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1. In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Botler et al. [2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC.
The concept of $m$-free graphs is particularly relevant to this paper, as it closely aligns with the work of Daamouch [7]. A directed graph $D$ is $m$-free if it contains no directed cycles of length at most $m$. In his paper, Daamouch introduces induced subgraphs and explores their role in the degree properties of vertices—an idea that will also play a significant role in the discussions and results presented here.
This paper builds on these foundations, generalizing an approach initially developed for the minimum degree 7 case [9]. Our key contribution is the introduction of a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers. This structure enables the construction of a decreasing sequence of subsets of nodes, forming the basis of our algorithmic proof of the SSNC. Our approach not only proves the conjecture but also identifies the specific node(s) whose out-degree doubles in the graph’s square.
Beyond its theoretical significance, the SSNC has practical applications in domains like social networks, epidemiology, and algorithm design. For instance, in the context of disease control, the conjecture helps identify individuals whose influence in their second-degree connections surpasses their direct contacts, offering insights into the spread of infections.
Unlike traditional proof-theoretic approaches, this work emphasizes the development of an efficient algorithm to locate the target node. We also analyze the complexity of our algorithm, demonstrating its practical feasibility for large-scale problems. The paper includes extensive supporting materials, such as detailed proofs, examples, and an interactive website to visualize key ideas, ensuring accessibility and transparency.
The remainder of this paper is structured as follows: Section 2 introduces common terminology used in graph theory. Section 3 outlines our approach to the problem, while Section 4 provides illustrative examples. Section 5 focuses on conditions under which a node’s out-degree may fail to double, and Section 6 examines the role of back arcs in oriented graphs. Section 7 explores the conditions that lead to doubling of interior degrees, examining their significance in relation to the SSNC. It introduces a divide-and-conquer approach, where the graph is partitioned, and each subproblem is solved independently before being integrated into the broader structure. Section 8 explores density constraints in containers. Section 9 investigates the impact of transitive triangles and Seymour diamonds on our algorithm. Finally, Section 10 presents the full Decreasing Sequence Property algorithm and its proof of the conjecture.
@@ 2 Graph Theory Terminology
Now that we have two different conjectures, we can investigate how they relate to one another. Conjecture 1.1 relates the notion of a node’s degree doubling to its neighborhoods. Similarly, conjecture 2.1 relates this same concept to the square graph. The defining characteristics of these conjectures are the relationship between a node’s out-degree at least doubling, and its second neighborhood being greater than or equal to the first. By the definition of an arc in the square graph, we can see that an arc is defined in this square graph only for paths of length two in the original graph. This is why these two conjectures are seen as identical.
We will focus our attention on Conjecture 1.1 and Conjecture 2.1. Conjecture 1.1 is easier to work with initially and understand why a node’s degree will double in the square graph. Concentrating on neighborhood analysis yielded numerous insights and features that aided in the development of our algorithm.
Next, we turn our attention to the term degree-doubling nodes. By this term, we aim to get to the heart of conjecture 1.1. That conjecture is looking for a node whose degree at least doubles. In order to do this, we examine the node’s first and second neighborhoods. Then identify a node whose second neighborhood is at least as large as the first, as required by the conjecture. A similar definition, a Seymour vertex, fulfills conjecture 2.1 - it’s out-degree in the squared graph is at least double its original out-degree. We could equivalently search for a Seymour vertex or a degree-doubling node.
@@ 3 Approach and Motivation
From this foundation, we can now partition the graph into $d_{v_{0}}+1$ subgraphs corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$. This partitioning enables us to leverage the divide-and-conquer paradigm. No nodes or arcs are being removed to form these partitions. This is a conceptual idea. Instead, think of it as putting a box or circle around the nodes and arcs. Now the nodes inside the box can work with the other nodes inside the box and solve their problem independent of the other partitions. These partitions are still connected to one another and will have to reconnect to solve the global problem. However, thinking of them in this way allows us to use their independent nature.
This represents a stark contrast to most previous approaches to the SSNC. Many of these previous approaches were conceptual or theoretical. Papers would mention the existence of a node whose degree doubled. The notion, though, of a path to that node was not part of the paper. Secondly, most of these papers focused on local search techniques that started at or near the desired degree-doubling node. That presents a problem because oriented graphs can be varying in nature. This is the first global search technique on the SSNC problem.
Complexity enters the picture when discussing global search algorithms. This raises concerns about how well they will handle larger graphs in comparison to smaller ones. This should not be a problem, as we will see in the following sections. The slowest part of this algorithm is building the data structure. That will require partitioning the nodes and edges of the graph, in a worst-case complexity of $O(V+E)$.
When the graph is partitioned, even larger graphs, often with higher value minimum degree nodes and thus more partitions, are able to be broken into subproblems. These subproblems are then able to be solved independent of one another. These subproblems are each their own instances of the SSNC, just a localized version. What the partition is doing is breaking a node’s individual degree into two parts—the degree within the container and the degree to the next container. The degree within the container is the subproblem that will be solved separately, and the degree to the next container is how they connect back to the global problem.
We still need to show that this is a formidable foundation. We have laid the groundwork for our algorithm to tackle the oriented graphs. With a minimum degree node, we can partition this graph into independent subproblems. This is now ready for a global search algorithm.
@@ 4 Initial Lemmas
These first lemmas will seek to place a lower bound on the out-degree of a minimum out-degree node in a counterexample. For this section, unless otherwise noted, assume that $v_{0}$ is a node with minimum out-degree in our oriented graph $G$. The work in this section agrees with what was already done by [11] and is not intended as original content. However, it does provide insight into our reasoning.
@@
Minimum Out-Degree $<$ 3 Suppose that a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double.
@@
Case 1: Assume that $d^{+}(v_{0})=0$. Then $v_{0}$ has no outgoing edges in $G$. Then $v_{0}$’s first and second neighborhoods are of size 0. Hence, its out-degree trivially satisfies the condition of at least doubling.
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. We know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. This node $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$. This means that $(v_{1},v_{2})\in G$. Hence $v_{2}$ is a second neighbor of $v_{0}$, thus making it at least as large as $v_{0}$’s first neighborhood.
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. We know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. Consider the induced subgraph $G[v_{1},v_{2}]$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. The node $v_{0}$ cannot be one of these nodes. Hence $v_{0}$ will have at least two second neighbors - $v_{3}$ and $v_{4}$, which makes its second neighborhood at least as large as its first. Thus, its out-degree will double in $G^{2}$.
Therefore, in all cases, $v_{0}$’s out-degree at least doubles.
Lemma 4.1 Minimum Out-Degree $<$ 3 begins to tell us about the structure of oriented graphs. We are learning that, to construct our counterexample, we need the minimum degree of our graph to be at least 3. In each of the cases of the lemma, the minimum degree node has few very out-neighbors, which couldn’t stop the increase that was coming from the influx of second neighbors.
@@
Minimum Out-Degree 3 with Neighbor Out-Degree 0 Suppose a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3. Suppose also that at least one of $v_{0}$’s neighbors has out-degree 0. Then the out-degree of $v_{0}$ will at least double in $G^{2}$.
@@
Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, and $v_{3}\in V(G)$. Consider the induced subgraph $G[v_{1},v_{2},v_{3}]$.
Now, assume that $v_{1}$ has an out-degree of 0 within $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to 3 nodes outside of $\{v_{1},v_{2},v_{3}\}$.
We know that $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has at least 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$. This means that $v_{0}$ has at least three second neighbors. As a result, $v_{0}$ has at least as many second neighbors as first neighbors, and equivalently, $v_{0}$’s out-degree doubles in $G^{2}$. ∎
@@
Figure 3: This is a representation of the node $v_{0}$ in $G$, $v_{0}$ being a minimum degree node, connecting to three nodes, $v_{1}$, $v_{2}$, and $v_{3}$. Each of these nodes ($v_{1}$, $v_{2}$, and $v_{3}$) connects to the nodes $v_{4}$ and $v_{5}$. The neighbor of $v_{4}$ is $v_{5}$, and the neighbors of $v_{5}$ are $v_{6}$, $v_{7}$, and $v_{8}$. While $v_{0}$ has degree 3, the neighbors of $v_{0}$ all have degree 1 in the subgraph $G[N^{+}(v_{0})]$, and they each have degree 2 outside $G[N^{+}(v_{0})]$. Thus, in $G^{2}$, the degree of $v_{1}$ will double. See lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1)
@@
Minimum Out-Degree 3 with Neighbors 1 Assume that in an oriented graph, the minimal out-degree node $v_{0}$ has an out-degree of 3. Additionally, suppose that in the induced subgraph of its neighbors, every neighbor has an out-degree of 1. Then the out-degree of $v_{0}$ or one of its neighbors will then at least double in $G^{2}$.
@@
Each node $v_{1},v_{2},v_{3}$ must connect to two nodes in $G-G[v_{1},v_{2},v_{3}]$. If the set of nodes to which $v_{1},v_{2},v_{3}$ connect in $G-G[v_{1},v_{2},v_{3}]$ contains more than 2 distinct nodes, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume that they all connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
This reduces the situation to Case 3 in Lemma 4.1 (Minimum Out-Degree $<$ 3), where $v_{1}$ acts similarly to how $v_{0}$ did in that case. There is at most one edge between $v_{4}$ and $v_{5}$. One of these nodes, say $v_{4}$, must connect to three additional nodes, $v_{6},v_{7}$, and $v_{8}$. As a result, $v_{4}$ will have three second neighbors, and that is at least as many as its first neighbors. This means that degree doubles in $G^{2}$. ∎
Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) provides a global search strategy, which is a question we shall revisit throughout this paper. It must prevent a node’s out-degree from doubling for that to function. In this instance, we observe that the degree of the starting node in the example $v_{0}$ is not doubled. We are compelled to look for the next node instead. In this case, the degree of the neighbor of $v_{0}$, $v_{1}$, doubles.
We did not require the ”right” starting node. Therefore, this is a crucial distinction. Not because of its possible proximity to the degree-doubling node. Instead, we chose a minimal degree node due to its ability to split the graph’s nodes.
We shall use double arrows to show that a collection of nodes connects to every node in the next induced subgraph as shown in example 4.2. Because all nodes in a shared induced subgraph relate to the same set of nodes, this is possible.